OpenStudy (anonymous):

Find the LCM. 6a3, 8a

OpenStudy (jojo4eva):

a

OpenStudy (jojo4eva):

were there any answer choices

OpenStudy (anonymous):

no 

OpenStudy (anonymous):

\6a^3 = 2*3*a*a*a 8a= 2*2*2*a LCM=2a

OpenStudy (anonymous):

LCM = 3*a*a*a*2*2*2=24a^3